# SOP-Bench: Complex Industrial SOPs for Evaluating LLM Agents

### v1

@@
Abstract
Large Language Models (LLMs) have demonstrated impressive general-purpose reasoning, language understanding and problem-solving abilities. However, they continue to fall short in executing complex, long-horizon workflows that require strict adherence to Standard Operating Procedures (SOPs), a critical requirement for real-world industrial automation. Despite this, there is a lack of publicly available benchmarks that capture the complexity, structure and domain-specific nuances of industrial SOPs. To address this gap, we present three main contributions. First, we introduce a novel synthetic data generation framework for creating realistic, industry-grade SOPs that rigorously test the planning, reasoning, and tool-use capabilities of LLM-based agents. Second, using this framework, we develop SOP-Bench, a benchmark comprising over 1,800 tasks across 10 industrial domains, each coupled with APIs, tool interfaces, and human-validated test cases. Third, we evaluate two prominent agent architectures: Function-Calling Agent and ReAct Agent, on SOP-Bench, and observe average task success rates of only 27% and 48%, respectively. Remarkably, in tasks where the tools registry is much larger than what is needed for task completion, the agents invoke an incorrect tool nearly 100% of the time. Our findings underscore a substantial gap between current LLM’s agentic capabilities and the demands of real-world SOPs’ successful automations.
Furthermore, performance varies considerably by task and domain, which highlights the need for domain-specific benchmarking and careful architectural selection prior to agent deployment. SOP-Bench is publicly available at http://sop-bench.s3-website-us-west-2.amazonaws.com/. We are also releasing all the prompts that underpin the complete data generation framework to support creation of new synthetic SOPs for domain-specific evaluation. We invite the community to extend SOP-Bench by contributing the SOPs from their own industrial domains.
1 Introduction
Standard Operating Procedures (SOPs) are essential to consistent and reliable operations across industries, encoding domain expertise and compliance rules into structured workflows [4, 5]. With recent advances in large language models (LLMs), there is growing interest in automating SOP execution using LLM-based agents [25, 30, 2]. However, deploying such agents in production poses challenges that are not captured by existing benchmarks like Gorilla [16], API-Bank [9], ComplexBench [27].
Criteria SLUHN-SOP UNTHS-SOP
@@ 1 Introduction
Instruction ambiguity Low Moderate
Complexity of steps High Low to moderate
Implicit knowledge needed High Low to moderate
@@
Table 1: Highlighting the diveristy of SOPs in industry, assessed by OpenAI’s GPT-4-turbo [15] and validated by human associates with extensive cross-domain SOP review experience.
@@
These challenges are particularly acute in knowledge operations like verification services, support ticket triaging, and customer support, where SOPs involve ambiguous language, implicit domain knowledge, and complex decision trees [21]. For example, consider the two patient registration SOPs in appendix E, one from St. Luke’s University Health Network (SLUHN) [^1] https://www.slhn.org/-/media/slhn/Research/File/PDF/SOPs-and-Policies/SOP-302Patient-Registration-and-Ongoing-Subject-MgmtV30CLEAN71216.ashx, and one from University of North Texas Health Science Center (UNTHS)  [^2] https://www.unthsc.edu/wp-content/uploads/sites/23/Patient_Registration.pdf. Table 1 shows how different two SOPs of the same domain can be.
Generally speaking, industrial knowledge operations SOPs, spanning domains like customer service, content moderation, and data annotation, present challenges distinct from traditional process automation tasks [20]. Unlike manufacturing processes, these operations require complex information synthesis and context-dependent decisions in ambiguous situations, with content moderators and annotators navigating evolving guidelines and subjective classifications [21]. These operations frequently encounter edge cases that challenge policy interpretations [6], making them particularly difficult for existing LLM-based automation approaches that assume well-defined success criteria and stable procedures. See figure 1 for an example of an instruction that can be easier for humans to follow but could be ambiguous for LLMs.
@@
While recent advances in LLMs have enabled sophisticated instruction following and tool use [26, 30], automating SOP execution poses unique challenges that existing benchmarks fail to capture. Current evaluation frameworks primarily focus on isolated capabilities—tool use [19], planning [30], or instruction following [26], using clean, synthetic prompts that sidestep the messiness of real-world procedures.
@@
Figure 1: Example of instructions that can be simple for humans but confusing for LLMs: In this excerpt of page 2 of the UNHTS-SOP (appendix E) we can see that steps 4 and 6 both have instructions on verifying the insurance, but does not have explicit information on how to verify. Moreover, it is not clear why there are two steps of insurance verifications. For human associates with implicit knowledge of the domains this might not be as ambiguous as for LLMs. In our experience, this is a common occurrence in industrial SOPs.
To address the gap, we introduce SOP-Bench, a synthetic yet realistic benchmark designed to evaluate LLM agents on SOP-style tasks. SOP-Bench spans 1,800+ tasks across ten domains - including seller appeals triage, and hazardous goods classification, and more. See appendix B for details. The SOPs are paired with synthetic APIs, executable tools, and human-validated test cases. SOP-Bench is designed to assess how agents handle complex control flows, edge cases, and interdependent steps characteristic of industrial SOPs.
We make three contributions: (1) A framework for generating realistic, domain-rich SOP datasets; (2) A benchmark that mirrors industrial SOP complexity while supporting agent evaluation; and (3) Empirical evidence showing that some of the most popular agent paradigms, Function-Calling [13] and ReAct [30], achieve average task success rates of only 27% and 48% respectively, with wide variation across task types, while achieving higher than 90% on tasks in ToolBench [28].
Overall, our goal is to establish a standardized benchmark for evaluating AI agents’ readiness for SOP execution. We encourage the community to use SOP-Bench to exhaustively test different agent paradigms based on different pre-trained models. We are also releasing all prompts and the full data generation framework to enable users to create new synthetic SOPs for domain-specific evaluation and contribute back to SOP-Bench, so that the whole community can benefit from an increasingly larger and more diverse evaluation set.
2 Related Work
@@
LLM Agents for Planning and Tool Use. LLMs have enabled the emergence of AI agents capable of task planning, reasoning, and tool use. Architectures such as ReAct [30], AutoGPT [29], and LangChain [2] have demonstrated autonomous decision-making via multi-step reasoning and API calls. ToolChain [31] and ToolLLM [18] extend this by integrating dynamic tool execution. However, these systems are mostly evaluated in highly simplistic settings, limiting their applicability to real-world operational workflows.
Benchmark
Inst.
Follow
Multi
Step
Tool
Use
Real
Ambig.
Error
Hand.
Work
Flow
Ind.
Ctx.
AlpacaEval [7] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
FollowEval [8] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
ComplexBench [27] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
InFoBench [17] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$
Gorilla [16] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
API-Bank [9] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
ToolBench [19] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
AgentBench [12] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
PlanBench [23] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
ALFWorld [22] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
SOP-Bench [Ours] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 2: Comparison of capabilities across different benchmarks. (Column abbreviations: Inst. Follow = Instruction Following (basic ability to follow instructions); Multi Step = Multi-step Tasks (handling sequential, dependent tasks); Tool Use = Tool/API Integration (ability to use external tools and APIs); Real Ambig. = Real-world Ambiguity (handling unclear or incomplete instructions); Error Hand. = Error Handling (ability to handle and recover from errors); Work Flow = Workflow Dependencies (managing inter-task dependencies); Ind. Ctx. = Industrial Context (relevance to real-world industrial settings)).
@@
Instruction Following and Benchmarking. Instruction-following ability has been benchmarked through datasets like SUPER-NATURAL-INSTRUCTIONS [26], AlpacaEval [10], and FollowEval [8]. More recent multi-step benchmarks like ComplexBench [27] and InFoBench [17] explore instruction complexity, but typically use machine-formatted inputs that omit the ambiguity and variability of human-authored SOPs.
@@
Tool Use and API-Centric Benchmarks. Benchmarks like Gorilla [16], API-Bank [9], BENCHAGENTS  [1] and ToolBench [19] assess tool invocation and API usage capabilities. However, these focus on isolated tool interactions with minimal procedural context. In contrast, SOP execution involves coordinated tool use across dependent steps, often requiring error handling and state tracking.
@@
Agent Evaluation Frameworks. AgentBench [12] and PlanBench [23] offer general frameworks to evaluate LLM agents on planning and tool use. Yet, their task settings are narrow in scope and lack the procedural structure, conditional logic, and real-world ambiguity typical of industrial SOPs. Similarly, embodied agents are evaluated in ALFWorld [22] and BabyAI [3], but those environments differ fundamentally from workflow automation tasks.
SOP Automation and Business Process Modeling. Traditional business process automation relies on rule-based systems and formal process modeling languages [5, 11, 24]. While effective, these require manual formalization of procedures, making them brittle and labor-intensive. Prior work on SOP formalization in organizational contexts [4] has not yet translated into LLM-native solutions.
@@
How SOP-Bench Differs SOP-Bench addresses these gaps by offering (1) realistic SOP-style instructions, (2) diverse domain coverage, (3) structured APIs for execution, and (4) evaluation protocols grounded in real-world complexity. Compared to benchmarks like Gorilla and ToolBench that isolate API usage, SOP-Bench evaluates agents on full workflows with interdependencies, ambiguity, and error handling requirements. It provides a more comprehensive testbed for assessing agent robustness in enterprise automation settings.
3 Generating Industry-grade SOPs
At the core of our work lies the SOP-Bench benchmark generation workflow, a novel LLM prompting pipeline that leverages Anthropic’s Claude 3.5 Sonnet v2 model in conjunction with human oversight to develop industrial-grade SOPs across a diverse spectrum of domains that mimics real-world procedures (Figure.  2). The pipeline employs a hierarchical prompting strategy with six specialized templates, that produces SOPs which capture the nuances of human-authored procedures such as industry-specific jargon, interdependent logic, implicit domain knowledge, ambiguous instructions, and cross-referenced content. The framework enables agent builders to systematically create domain specific benchmarks with appropriate datasets and tools needed to evaluate the LLM agent’s performance on sequential reasoning, tool use and constraint adherence in procedural knowledge tasks. The benchmark generation workflow has two major steps: in the first stage, all the components of the benchmark such as SOP, tool specifications and code are generated without introducing any unnecessary challenges, ensuring that the core procedural logic is clearly expressed and coherent. In the second stage, we selectively introduce complexity in order to make the SOP as realistic as possible to real-world conditions by modifying the SOP text to include additional constraints, dependencies, or ambiguities such as incomplete information as well as adding redundant tools to challenge the LLM agent’s reasoning and decision-making in tool selection.
Figure 2: SOP-Bench Synthetic Benchmark Generation Workflow using LLMs and Human oversight. ** denotes post-generation complexity introduction in respective components.
3.1 SOP-Bench Benchmark Generation Workflow
The benchmark generation process uses two inputs from the user namely: <Business Task, Task Context>. These two inputs form the critical backbone of the benchmark generation process. A detailed example is provided in  5
Business Task: A one line description of the business task that clearly articulates the topic and purpose of the SOP to be created.
@@ 3.1 SOP-Bench Benchmark Generation Workflow
Task Context: A brief description of the different procedural steps required in the generated SOP in order to process a given user request successfully. For the sample patient intake SOP, this includes steps such as validating patient information, insurance details, previous medical history and assessing risk factors.
Here is an example of business task and task context which includes high-level business objectives that is used to guide SOP generation.
  • Business Task
    • SOP outlining the steps for registering new patients
  • Task Context
    • Describe the purpose and scope of the patient intake process specific to new patients
    • Validate insurance information and eligibility
    • Validate prescription insurance
    • ... Continued ...
3.1.1 Dataset Schema Generation
Instead of directly prompting the LLM with <Business Task, Task Context> to generate SOPs, we first generate a structured dataset schema using one-shot prompting (see Appendix C.2). This schema outlines real-world task scenarios, specifying input parameters, intermediate logic, decision points, compliance checks, and expected outcomes. This step is crucial for minimizing hallucinations and correcting structural inconsistencies, thereby establishing a reliable foundation for SOP and tool generation. Acting as a semantic scaffold, the schema embeds critical operational context that guides and constrains subsequent SOP generation, ensuring outputs are both relevant and executable. With human (domain expert) oversight, we validate that the schema captures essential task components, including implicit knowledge and domain-specific constraints. As illustrated in the following example, When the data schema is not referenced, the SOP instructions remain vague whereas adding the dataset schema ensures instructions are precise and consistent. For example,
  • Schema Entry

### v2

@@
Abstract.
LLM-based agents struggle to execute complex, multi-step Standard Operating Procedures (SOPs) that are fundamental to industrial automation. Existing benchmarks fail to capture the procedural complexity and tool orchestration demands of real-world workflows. We introduce SOP-Bench, a benchmark of 2,000+ tasks from human expert-authored SOPs across 12 business domains (healthcare, logistics, finance, content moderation, etc.). Using a human-AI collaborative framework, experts crafted authentic SOPs while AI generated artifacts (tools, APIs, datasets), all human-validated, yielding realistic tasks with executable interfaces and ground-truth outputs.
SOP-Bench serves as a research enabler for systematically investigating agent architectures, model capabilities, and deployment considerations across diverse procedural tasks. We demonstrate its utility through illustrative experiments with a subset of frontier models across Function-Calling (FC) and ReAct agents, revealing critical insights. For example, (1) newer models do not guarantee better performance - Claude 4 family outperforms Claude 4.5 family on ReAct tasks (Claude 4 Opus: 72.4% vs. Claude 4.5 Sonnet: 63.3% task success rate), demonstrating that production upgrades require validation; (2) no single model-agent combination dominates: best performances range from 57% to 100% depending on domain. These examples illustrate how SOP-Bench enables isolating and studying specific dimensions of agent performance without costly production experiments. Our goal is not to rank model capabilities or build optimal agents, but to provide a rigorous evaluation framework that enables the researchers and practitioners to systematically investigate agent design choices, model selection, and deployment strategies. We release the benchmark at https://github.com/amazon-science/sop-bench.
copyright: none
Figure 1. SOP-Bench evaluation overview. Realistic business process SOPs authored by human experts across diverse domains are converted into executable task instances with structured tool/API interfaces and ground-truth outputs. LLM agents execute tasks via reproducible tool interactions, producing execution trajectories evaluated using grounded, outcome-aware metrics (ECR, C-TSR, TSR).
1. Introduction
Standard Operating Procedures (SOPs) are central to reliable operations across industries, encoding domain expertise, compliance rules, and decision logic into structured workflows (De Treville et al., 2009; Dumas et al., 2018). Recent advances in large language models (LLMs) have sparked growing interest in automating SOP execution using LLM-based agents (Wang et al., 2023; Yao et al., 2023; Chase, 2022; Gschwind et al., 2025). However, deploying such agents in production exposes challenges that are not captured by existing benchmarks such as Gorilla (Patil et al., 2023), API-Bank (Li et al., 2023a), or ComplexBench (Wen et al., 2024).
Table 1. Diversity of SOPs in industry, assessed by GPT-4-turbo (OpenAI, 2024) and validated by human associates with extensive cross-domain SOP review experience.
Criteria SLUHN UNTHS
@@ 1. Introduction
Ambiguity Low Moderate
Step complexity High Low-Mod
Implicit knowledge High Low-Mod
@@
These challenges are particularly acute in knowledge operations like verification services, support ticket triaging, and customer support, where SOPs involve ambiguous language, implicit domain knowledge, and complex decision trees (Roberts, 2019). For example, consider the two patient registration SOPs in appendix E, one from St. Luke’s University Health Network (SLUHN) [^1] https://www.slhn.org/-/media/slhn/Research/File/PDF/SOPs-and-Policies/SOP-302Patient-Registration-and-Ongoing-Subject-MgmtV30CLEAN71216.ashx, and one from University of North Texas Health Science Center (UNTHS)  [^2] https://www.unthsc.edu/wp-content/uploads/sites/23/Patient_Registration.pdf. Table 1 shows how different two SOPs of the same domain can be.
Such business knowledge operations differ fundamentally from traditional process automation tasks (Ribeiro et al., 2020). They require contextual interpretation, multi-step information synthesis, and judgment under ambiguity, often with evolving guidelines and subjective classifications (Roberts, 2019; Gordon et al., 2022). Figure 2 shows a representative example: instructions that are straightforward for human operators with implicit domain knowledge but ambiguous for LLM agents. Existing evaluation frameworks typically isolate individual capabilities—such as tool use (Huang et al., 2024), planning (Yao et al., 2023), or instruction following (Wang et al., 2022)—using clean, synthetic prompts that fail to reflect the messiness of real-world SOPs.
@@
While recent advances in LLMs have enabled sophisticated instruction following and tool use (Wang et al., 2022; Yao et al., 2023), automating SOP execution poses unique challenges that existing benchmarks fail to capture. Current evaluation frameworks primarily focus on isolated capabilities—tool use (Huang et al., 2024), planning (Yao et al., 2023), or instruction following (Wang et al., 2022), using clean, synthetic prompts that sidestep the messiness of real-world procedures.
@@
Figure 2. Example of instructions that can be simple for humans but confusing for LLMs. In this excerpt of page 2 of the UNHTS-SOP (appendix E), we can see that steps 4 and 6 both have instructions on verifying the insurance, but does not have explicit information on how to verify it. Moreover, it is not clear why there are two steps of insurance verifications. For human associates with implicit knowledge of the domains this might not be as ambiguous, while our experiments indicate that it is for LLMs. In our experience, this is a common occurrence in business process SOPs.
Screenshot of a section of the UNHTS SOP showing numbered steps. Two different steps mention verifying insurance, but neither specifies how to verify or why verification is repeated.
To address this gap, we introduce SOP-Bench, a comprehensive benchmark for evaluating LLM agents on realistic business SOP execution. SOP-Bench is constructed from human expert-authored SOPs spanning twelve diverse domains, including healthcare intake, hazardous goods classification, customer service, content moderation, and financial compliance. Industry experts authored authentic procedures reflecting real-world ambiguity, branching logic, and implicit knowledge requirements. To enable reproducible evaluation, we adopt a human–AI collaborative workflow in which AI models generate supporting artifacts (mock tools, APIs, and test datasets) that are subsequently validated by human experts. This process yields over 2,000 executable tasks with structured tool interfaces and grounded outputs that closely mirror production environments.
We make three primary contributions: (1) Human–AI collaborative benchmark construction framework: We introduce a scalable methodology where industry experts author authentic SOPs while AI assists in generating executable artifacts, all human-validated for correctness and realism. (2) A comprehensive benchmark for systematic agent evaluation: SOP-Bench provides 2,000+ tasks across 12 domains with varying procedural complexity, document length, branching logic, and tool usage, enabling controlled study of agent architectures and foundation models without costly production experimentation. (3) Empirical insights via illustrative experiments: Through experiments and ablations on representative frontier models, we demonstrate how SOP-Bench reveals non-obvious trade-offs in architecture choice, reasoning cost, and domain generalization.
Our goal is not merely to rank models, but to establish SOP-Bench as an evaluation substrate for studying agent design, model selection, and deployment strategies under realistic procedural constraints. We release the full benchmark, baseline agents, and evaluation framework at https://github.com/amazon-science/sop-bench, enabling the community to extend SOP-Bench with new domains and contribute back to a shared evaluation ecosystem.
Table 2. SOP-Bench uniquely combines all seven capabilities essential for real-world industrial workflow automation, addressing critical gaps in existing agent benchmarks.
Benchmark IF MS Tool Amb. Err. Dep. Ind.
AlpacaEval
FollowEval
ComplexBench
InFoBench
Gorilla
API-Bank
AgentBench
PlanBench
ALFWorld
SOP-Maze
SOP-Bench
Legend: IF = Instruction Following; MS = Multi-step Tasks; Tool = Tool/API Integration; Amb. = Real-world Ambiguity; Err. = Error Handling; Dep. = Workflow Dependencies; Ind. = Industrial Context.
2. Related Work
@@
LLM Agents for Planning and Tool Use. LLMs have enabled the emergence of AI agents capable of task planning, reasoning, and tool use. Architectures such as ReAct (Yao et al., 2023), AutoGPT (Yang et al., 2023), and LangChain (Chase, 2022) have demonstrated autonomous decision-making via multi-step reasoning and API calls. ToolChain (Zhuang et al., 2023) and ToolLLM (Qin et al., 2023) extend this by integrating dynamic tool execution. However, these systems are mostly evaluated in highly simplistic settings, limiting their applicability to real-world operational workflows.
@@
Instruction Following and Benchmarking. Instruction following ability has been benchmarked through datasets like SUPER-NATURAL-INSTRUCTIONS (Wang et al., 2022), AlpacaEval (Li et al., 2023b), and FollowEval (Jing et al., 2023). More recent multi-step benchmarks like ComplexBench (Wen et al., 2024) and InFoBench (Qin et al., 2024) explore instruction complexity, but typically use machine-formatted inputs that omit the ambiguity and variability of human-authored SOPs.
@@
Tool Use and API-Centric Benchmarks. Benchmarks like Gorilla (Patil et al., 2023), API-Bank (Li et al., 2023a) and BENCHAGENTS  (Butt et al., 2024) assess tool invocation and API usage capabilities. However, these focus on isolated tool interactions with minimal procedural context. In contrast, SOP execution involves coordinated tool use across dependent steps, often requiring error handling and state tracking.
@@
Agent Evaluation Frameworks. AgentBench (Liu et al., 2023) and PlanBench (Valmeekam et al., 2023) offer general frameworks to evaluate LLM agents on planning and tool use. Yet, their task settings are narrow in scope and lack the procedural structure, conditional logic, and real-world ambiguity typical of industrial SOPs. Similarly, embodied agents are evaluated in ALFWorld (Shridhar et al., 2020) and BabyAI (Chevalier-Boisvert et al., 2018), but those environments (e.g., a kitchen) differ fundamentally from workflow automation tasks.
SOP Automation and Business Process Modeling. Traditional business process automation relies on rule-based systems and formal process modeling languages (Dumas et al., 2018; Lindsay et al., 2003; Van Der Aalst et al., 2003). While effective, these require manual formalization of procedures, making them brittle and labor-intensive. Prior work on SOP formalization in organizational contexts (De Treville et al., 2009) has not yet translated into LLM-native solutions. Recently, (Gschwind et al., 2025) proposes a novel LLM Classifier-Augmented Generation (CAG) approach that translates procedures described in natural language into executable workflows. However, their target descriptions are rather short and limited to the domain of data integration. Most importantly, the dataset used to evaluate the approach is not publicly available. Recently, SOP-Maze (Wang et al., 2025) has published business operations SOPs, but they lack the tools or ground truth data for agent evaluation.
@@
How SOP-Bench Differs SOP-Bench addresses these gaps by offering (1) realistic SOP-style instructions, (2) diverse domain coverage, (3) structured APIs for execution, and (4) evaluation protocols grounded in real-world complexity. Compared to benchmarks like Gorilla that isolate API usage, SOP-Bench evaluates agents on full workflows with inter-dependencies, ambiguity, and error handling requirements. It provides a more comprehensive testbed for assessing agent robustness in enterprise automation settings.
3. Generating Industry-grade SOPs
SOP-Bench is built on a novel human-AI collaborative framework where human domain experts author authentic SOPs while AI models generate supporting artifacts, all subsequently human-validated. This approach ensures procedural authenticity while maintaining reproducibility and scalability. The workflow employs a hierarchical prompting strategy using Anthropic’s Claude 3.5 Sonnet v2 to refine human-authored SOPs for consistency and completeness, and to generate associated artifacts (tools, APIs, datasets) that capture real-world nuances: industry jargon, interdependent logic, implicit domain knowledge, ambiguous instructions, and cross-referenced content (Figure 3).
Figure 3. SOP-Bench Generation Workflow: Human experts author SOPs and validate all artifacts; AI generates tools, APIs, and datasets. ** denotes post-generation complexity introduction.
Diagram illustrating the SOP-Bench generation workflow. Human experts create and validate SOPs, while an AI system generates tools, APIs, and datasets. Certain stages are annotated to indicate additional complexity introduced after generation.
3.1. Human-AI Collaborative Workflow
Human Expert Input: Domain experts provide two critical inputs that define the benchmark scope:
Business Task: The first version of an industrial use-case SOP
@@ 3.1. Human-AI Collaborative Workflow
Task Context: Procedural steps required to process requests successfully. For patient intake, this includes validating patient information, insurance details, medical history, and risk assessment.
  • Business Task: SOP outlining steps for registering new patients
  • Task Context: Validate insurance eligibility, prescription coverage, ...
3.1.1. Dataset Schema Generation (LLM-Assisted, Human-Validated)
The first step is to create a structured dataset schema via one-shot prompting (Appendix C.2). This schema specifies input parameters, decision points, compliance checks, and expected outcomes—acting as a semantic scaffold that minimizes hallucinations and ensures consistency. Human domain experts validate that the schema captures essential task components, including implicit knowledge and domain constraints. It uses the SOP and task context to generate the dataset schema. For example: